Solve the equation.
6x^(-1/2) - 13x^(-1/4) +6 = 0


** Someone tried to help me with this saying to set w = x^(-1/4) but thats wrong when you take w^2 it does not give you (1/2) it gives you (1/8)

Answer by palanisamy(496)

The given equation is
6x^(-1/2) - 13x^(-1/4) +6 = 0
Put x^(-1/4) = y. Then x^(-1/2) = y^2
So, we get
6y^2-13y+6=0
6y^2-9y-4y+6=0
3y(2y-3)-2(2y-3)=0
(2y-3)(3y-2)=0
2y-3=0 or 3y-2 = 0
y = 3/2 or 2/3
x^(-1/4)= 3/2 or x^(-1/4)= 2/3
x = (3/2)^(-4) or x = (2/3)^(-4)
x = (2/3)^4 or x = (3/2)^4
x=16/81 or 81/16